# ALoRA: Allocating Low-Rank Adaptation for Fine-tuning Large Language Models

### v1

@@ 3.3 Our novel AB-LoRA method
1 Train super-network $M$ on the training set $D_{train}$ for $K_{1}$ epochs; for $n=0$; $n<N_{A}$ do
@@ 3             Calculate the importance score $\text{IS}(r_{m,i})$ on $B_{val}$;
4      Prune $n_{0}$ LoRA ranks with lowest importance scores; if there are modules not pruned then
@@ Appendix I Case studies of Instruction tuning
Query
Edit the following paragraph to correct any grammatical errors: She didn’t remembre where is her purse, so I thinks its in the car but he’s say it’s on kitchen table but he are not sure, and then they asked me to looking for it, she’s say, "Can you?", and I responds with, "Maybe, but ain’t no sure," and he not heard me, and, "What?", he asks, "Did you found it?".
@@
Responses
She didn’t remember where her purse was, so I thought it was in the car, but he said it was on the kitchen table, but he wasn’t sure, and then they asked me to look for it, she said, "Can you?" and I responded with, "Maybe, but I’m not sure," and he didn’t hear me, and then he asked, "Did you find it?"
@@
Assistant 1’s response effectively corrects all the grammatical errors in the original paragraph. The assistant has correctly changed "remembre" to "remember," rearranged the clause order from "where is her purse" to "where her purse was," and adjusted verb forms for proper agreement (e.g., changing "thinks" to "thought"). Quotation marks and commas are properly used, making direct speech clear and punctuated correctly. The conjunctions are better placed, contributing to a more natural flow of sentences. Additionally, all instances of colloquial or incorrect language such as “ain’t no sure” have been revised into standard English ("I’m not sure"). Overall, this response would be very helpful to human users as it provides a grammatically sound version of the original text.
@@
Responses
She didn’t remember where her purse was, so I thought it was in the car, but he said it was on the kitchen table, but he wasn’t sure. They asked me to look for it, and she said, "Can you?" I responded with, "Maybe, but I’m not sure." He didn’t hear me, and he asked, "What?" I said, "Did you find it?"
@@
Assistant 2 also does an excellent job correcting grammatical mistakes from the initial query. This assistant similarly fixes issues with verb tense consistency and subject-verb agreement (e.g., “he’s say” corrected to “he said”). However, there is a slight deviation at the end where Assistant 2 writes: “I said, ’Did you find it?’”. This seems like an attempt by Assistant 2 to make sense of a confusing part of dialogue in which speaker roles were unclear but results in altering the meaning slightly because it suggests that ‘I’ am asking if ‘he’ found it rather than him asking ‘me’. This could be misleading or confusing without additional context indicating who is speaking at that moment. Nevertheless, punctuation around quotations is accurate which aids readability.

### v2

@@ 3.3 Our novel AB-LoRA method
2
3Train super-network $M$ on the training set $D_{train}$ for $K_{1}$ epochs;
4for $n=0$; $n<N_{A}$ do
@@ 6             Calculate the importance score $\text{IS}(r_{m,i})$ on $B_{val}$;
7
8      Prune $n_{0}$ LoRA ranks with lowest importance scores;
9      if there are modules not pruned then
@@ Appendix I Case studies of Instruction tuning
Query Edit the following paragraph to correct any grammatical errors: She didn’t remembre where is her purse, so I thinks its in the car but he’s say it’s on kitchen table but he are not sure, and then they asked me to looking for it, she’s say, "Can you?", and I responds with, "Maybe, but ain’t no sure," and he not heard me, and, "What?", he asks, "Did you found it?".
@@
Responses She didn’t remember where her purse was, so I thought it was in the car, but he said it was on the kitchen table, but he wasn’t sure, and then they asked me to look for it, she said, "Can you?" and I responded with, "Maybe, but I’m not sure," and he didn’t hear me, and then he asked, "Did you find it?"
@@
Responses She didn’t remember where her purse was, so I thought it was in the car, but he said it was on the kitchen table, but he wasn’t sure. They asked me to look for it, and she said, "Can you?" I responded with, "Maybe, but I’m not sure." He didn’t hear me, and he asked, "What?" I said, "Did you find it?"